carmelle:

Pick a two-digit number greater than 25. Rewrite your two-digit number as a difference of two numbers. Show how to use the identity (x − y)^2 = x^2 − 2xy + y^2 to square your number without using a calculator. The number is 30.

carmelle:

@AZ can you help?

carmelle:

@darkknight can you help?

carmelle:

@snowflake0531 can u help :)

snowflake0531:

Well choose 2 numbers _-_ = 30

Florisalreadytaken:

the point of this is to find what the square of that number would be, without using a calculator ok, so you picked 30 lets rewrite that as a difference of 2 numbers -- so that would be \(35-5\) where \( x=35 \) \( y=5 \) substitute that into this \((x − y)^2 = x^2 − 2xy + y^2\)

Florisalreadytaken:

actually, whats the point of this?😂

Florisalreadytaken:

you would use that calculator to find \( 35^2 \) anyway

carmelle:

ok \[(35 - 5)^{2} = 35^2 - 2(35)(5) + y^2\]

carmelle:

5^2 *

carmelle:

@florisalreadytaken wrote:
you would use that calculator to find \( 35^2 \) anyway
fr smh

carmelle:

1225 - 350 + 25 is this correct?

Florisalreadytaken:

ok, sp by doing the whole thing in 1 go, it will be: \( (x − y)^2 = x^2 − 2xy + y^2 \) \( (35 - 5)^{2} = 35^2 - 2(35)(5) + 5^2 \) \( 900 =1225 -350 + 25 \) \( 900= 900\)

carmelle:

thank you sooo much :D

Florisalreadytaken:

you got it :D

AZ:

@florisalreadytaken wrote:
actually, whats the point of this?😂
@florisalreadytaken wrote:
you would use that calculator to find \( 35^2 \) anyway
Perhaps if you used simpler numbers like 30^2 = (40-10)^2 then you wouldn't have to use a calculator but the purpose of this assignment seems to be just teaching you the identity and not necessarily the most efficient way of finding the square of a number